OpenStudy (anonymous):

Evaluate the expression -(sqare root of) 0.81

OpenStudy (lgbasallote):

\[\huge -\sqrt{81}\] do you know what the square root of 81 is?

OpenStudy (anonymous):

nope

OpenStudy (lgbasallote):

you're not allowed to use a calculator?

OpenStudy (anonymous):

9

OpenStudy (lgbasallote):

good..so negative square root of 81 is just negative of that

OpenStudy (anonymous):

so its the same, just with .81 not 81

OpenStudy (lgbasallote):

- not .

OpenStudy (anonymous):

well i said negative of that so would it be positive?

OpenStudy (lgbasallote):

wait..i got confused

OpenStudy (anonymous):

I put a negative before that :)

OpenStudy (lgbasallote):

\[-\sqrt{81}\] this is just negative of \(\sqrt {81}\)

OpenStudy (lgbasallote):

that's what i was referring to

OpenStudy (lgbasallote):

take the square root of 81 then put a negative sign

OpenStudy (anonymous):

thx bai